FMARS-15 XO Report 26-07-2023

 In XO Report

Sol: 3
Summary Title: Red Planet Explorations
Author’s Name: D. Terry Trevino
Mission Status: All Nominal

Sol Activity Summary: Crew 15 embarked on another productive day on Mars today. We started
the morning with a hearty breakfast of cereal and pancakes before commencing preparations for
today’s Extravehicular Activity (EVA). Crew members gathered their gear and checked their
space suits thoroughly to ensure their functionality during the EVA.

EVA activities included a revisit to previously explored sites in the Haughton Crater on the
Canadian Crown land. The objective was to collect additional geological and biological samples
and survey the area to observe potential changes. The EVA team, consisting of Commander
Andrew Wheeler, Health and Safety Officer Caleb Pool, spent approximately 4 hours outside.
During this time, Andy Greco, dutifully served as a bear guard, maintaining security for the

Back at the habitat, we continued our efforts to improve habitation conditions. Olivia and I
worked in the science bay, continuing our research on fluorescent photography to detect lipids
and nano plastics in the local water supply. The goal is to ensure the safety and sustainability of
our water resources on Mars.

To promote crew well-being and reduce stress, we practiced meditation sessions twice today
while monitoring our heart rate variability using our Garmin watches. We are on track to achieve
our target of three meditation sessions per day, which will aid in maintaining psychological
health during the mission.

Look Ahead Plan: As we proceed to Sol 4, we have outlined several important tasks:

Fuel Line Repair: The engineering team will continue repairing the fuel line leak in Genny. This
is crucial for ensuring a stable power supply to the habitat.
EVA Preparation: We plan to conduct another EVA in the coming days. Crew members will
prepare the necessary equipment and thoroughly check the space suits and backpacks.
Habitat Maintenance: Our team will declutter the engineering rooms and address lighting issues
in the upper hab. Additionally, we will focus on locating and resolving the CO2 anomaly to
maintain a healthy living environment.
Research and Reporting: Science Officer, Olivia Drayson and I will continue our research on
fluorescent photography, and the results will be documented in the science report. The required
reports, including Sol Summary, Operations Report, Journalist Report, EVA Report, and EVA
Request, will be diligently prepared and submitted as scheduled.
Weather: Today, the weather on Mars was sunny and bright, with temperatures ranging around
10°C. The humidity remained low, and the wind speed increased mildly.

Crew Physical Status: The overall physical health of the crew remains nominal. As planned, we
will share our personal health numbers for reference and monitoring.

Support Requested: We thank Eleonore and Sam for their valuable assistance in identifying
pollutants and clarifying concerns related to TVOC and formaldehyde. As we continue to
address air quality concerns, we appreciate further support and advice on ventilating the habitat
effectively over the upcoming winter and preparing for future crew arrivals.

Mission Success and Mars Exploration: We are proud of our crew’s accomplishments thus far,
and our mission remains on track. Each day brings us closer to understanding and adapting to
this unique environment on Mars. Our exploration of the Simulated Red Planet continues to yield
valuable data and insights that will contribute to future missions and humanity’s long-term
presence on Mars.

End of Report

Start typing and press Enter to search